# Demystifying Domain-adaptive Post-training for Financial LLMs

### v1

@@ Abstract
Domain-adaptive post-training of large language models (LLMs) has emerged as a promising approach for specialized domains such as medicine and finance. However, significant challenges remain in identifying optimal adaptation criteria and training strategies across varying data and model configurations. To address these challenges, we introduce FinDaP, a systematic and fine-grained investigation into domain-adaptive post-training of LLMs for the finance domain. Our approach begins by identifying the core capabilities required for the target domain and designing a comprehensive evaluation suite aligned with these needs. We then analyze the effectiveness of key post-training stages, including continual pre-training, instruction tuning, and preference alignment. Building on these insights, we propose an effective training recipe centered on a novel preference data distillation method, which leverages process signals from a generative reward model. The resulting model, Llama-Fin, achieves state-of-the-art performance across a wide range of financial tasks. Our analysis also highlights how each post-training stage contributes to distinct capabilities, uncovering specific challenges and effective solutions, providing valuable insights for domain adaptation of LLMs.
Project page: https://github.com/SalesforceAIResearch/FinDap
@@ 1 Introduction
While large language models (LLMs) demonstrate strong generalization across diverse tasks, they often struggle in specialized domains such as finance, law, and healthcare. Consequently, domain-adaptive post-training of LLMs has garnered significant attention recently (Colombo et al., 2024b; Xie et al., 2024a). In the earlier days of language models, continual pre-training or CPT was the dominant strategy. This involved further training a pre-trained model on domain-specific plain text and then fine-tuning it for individual tasks (Gururangan et al., 2020; Ke et al., 2023). While prompt engineering with zero- or few-shot examples has emerged as a convenient way to adapt general-purpose LLMs to new tasks, recent methods focus on fine-tuning model weights to transform them into domain experts (Colombo et al., 2024a; Chen et al., 2023a; Li et al., 2023). [^1] This shift is understandable not only because training typically leads to better results on specialized tasks, but also because training has become more affordable due to the growing number of techniques that enable efficient training. For example, the training cost was over \$4M for GPT-3 (davinci), but it has recently dropped to around \$0.8M to train a GPT-3 on-par model like Phi3.5  (Maslej et al., 2024).
@@
Today, domain-adaptive post-training has evolved into a collection of techniques including CPT, instruction-tunning (IT) and preference alignment (PA). Building on this trend, this work focuses on adapting LLMs to specific domains through parameter training, which complements semi-parametric methods that leverage external knowledge, such as in retrieval-augmented generation or RAG (Lewis et al., 2020; Ke et al., 2024). Our focus is also different from general post-training, as the goal is not to develop another general-purpose frontier LLM but to create specialized, expert-level LLMs tailored to a specific domain.
@@
Despite the need and aforementioned advancements in domain-specific post-training, there is still no systematic study on what makes a good domain-specific LLM (see Table 2). In this paper, we wish to address the following research questions:
@@
Research Questions Given a pre-trained LLM with strong general capabilities (e.g., Llama3-8b-instruct), how to effectively adapt it to a target domain by post-training? What criteria are desirable for successful adaptation? What are effective training recipes with respect to data and model?
A common criterion for evaluating domain-specific LLMs is to test their performance on a set of domain-specific tasks (Bhatia et al., 2024; Xie et al., 2024b). However, this simplified approach can often misalign with our broader expectations for a domain-expert LLM. The absence of clear and comprehensive evaluation criteria has significantly hindered the development of robust training recipes, encompassing both data and model strategies. Furthermore, it remains uncertain whether simply using a larger LLM will yield better performance in a specific domain compared to a carefully post-trained, domain-adaptive LLM.
To address these questions methodically, we introduce FinDaP, a novel finance-specific post-training framework comprising a comprehensive evaluation framework, state-of-the-art model checkpoints and a training recipe. We use the finance domain as a case study to perform domain-adaptive post-training on an instruction-tuned LLM, Llama3-8b-instruct (LLaMA, 2024). Figure 1 outlines our approach. We begin by identifying a set of core capabilities that a domain-expert LLM should possess such as knowledge about concepts and tasks in the target domain and the associated reasoning processes (§2.1). This identification is critical as it informs the entire exploration process. Based on these expected capabilities, we develop an evaluation framework (§2.2) to establish clear performance goals and guide model improvement over a selection of development and unseen tasks.
@@
The training recipe involves two stages. In the first stage (§3.2 & §3.3), we conduct rigorous and principled experiments to determine the optimal data and model recipes for CPT and IT, resulting in the ‘CPT+IT’ checkpoint. Subsequently, we further train the model using curated on-policy preference data for PA. We propose a novel approach to construct preference data from the process signals given by a generative reward model. Our best performing recipe yields Llama-Fin that outperforms all considered baselines, including large open models at the 70B scale and proprietary models like GPT-4o, on unseen tasks that are similar to the training data. Even on unseen and novel tasks never encountered in training, Llama-Fin remains competitive and consistently outperforms its base model across all identified capabilities. In summary, our key contributions are:
  • A comprehensive guidance for domain-specific post-training, including identification of capabilities, evaluation, data and model recipe design.
  • A systematic exploration on each stage of domain-specific post-training, with an emphasis on the goals, challenges and effective approaches at each stage.
  • A novel preference alignment approach that constructs preference data using on-policy trajectories guided by outcome and process signals from a generative reward model.
  • A new state-of-the-art finance LLM (Llama-Fin) at the 8b parameter scale.
  • Extensive evaluation with an open-source leaderboard, checkpoints, data and model recipes.
Figure 1: An overview of FinDaP. Left: we first identify the core expected capabilities for the target domain and then curate texts and prompts for training and evaluation. Right: the top shows our training strategies. For each training stage, we use development set to select the best model. After training, we use unseen set to demonstrate the effectiveness of Llama-Fin.
2 FinDaP Setup
In FinDaP, we first identify the desired capabilities and design the corresponding evaluation measures, and then formulate the training method.
2.1 FinDaP Capabilities
We begin by illustrating the capabilities that are desirable for a domain-specific LLM. Specifically, we focus on the following core capabilities:

### v2

# Domain-adaptive Post-training of Large Language Models: An Empirical Study with Llama for Finance [Yifei: A more attractive/informative title?]

@@ Abstract
Domain-adaptive post-training of large language models (LLMs) has emerged as a promising approach for specialized domains such as medicine and finance. However, significant challenges remain in identifying optimal adaptation criteria and training strategies across varying data and model configurations. To address these challenges, we introduce FinDaP, a systematic and fine-grained investigation into domain-adaptive post-training of LLMs for the finance domain. Our approach consists of four key components: FinCap, which defines the core capabilities required for the target domain; FinRec, an effective training recipe that jointly optimizes continual pre-training and instruction-following, along with a novel preference data distillation method leveraging process signals from a generative reward model; FinTrain, a curated set of training datasets supporting FinRec; and FinEval, a comprehensive evaluation suite aligned with FinCap. The resulting model, Llama-Fin, achieves state-of-the-art performance across a wide range of financial tasks. Our analysis also highlights how each post-training stage contributes to distinct capabilities, uncovering specific challenges and effective solutions, providing valuable insights for domain adaptation of LLMs.
Demystifying Domain-adaptive Post-training for Financial LLMs
Zixuan Ke, Yifei Ming, Xuan-Phi Nguyen, Caiming Xiong and Shafiq Joty Salesforce AI Research {zixuan.ke,yifei.ming,xnguyen,cxiong,sjoty}@salesforce.com [Uncaptioned image] Project Page: https://github.com/SalesforceAIResearch/FinDAP [Uncaptioned image] Datasets: https://huggingface.co/datasets/Salesforce/FinEval
@@ 1 Introduction
While LLMs have demonstrated strong generalization across a variety of tasks, they often struggle to perform well in specialized domains such as finance and law. Consequently, domain-adaptive post-training of LLMs has garnered significant attention recently (Colombo et al., 2024a; Xie et al., 2024b). In the earlier days of language models, continual pre-training (CPT) was the dominant strategy. This involved further training a pre-trained model on domain-specific plain text and then fine-tuning it for individual tasks (Gururangan et al., 2020; Ke et al., 2023). With LLMs, the post-training focus has shifted to zero- and few-shot task generalization through methods such as instruction-tunning (IT) (aka. supervised fine-tuning or SFT) and preference alignment (PA). While prompt engineering of powerful general LLMs with zero- or few-shot examples has emerged as a convenient approach to adapting them to new tasks, to get the most optimal performance on a target domain, recent methods explore fine-tuning model wights to make them domain experts Chen et al. (2023b); Li et al. (2023); Colombo et al. (2024b).
@@
Building on this trend, this work focuses on adapting LLMs to specific domains through parameter training. It complements semi-parametric methods that leverage external knowledge, such as retrieval-augmented generation (Lewis et al., 2020; Ke et al., 2024). Our focus is also different from general post-training, as the goal is not to develop another general-purpose LLM but to create specialized, expert-level LLMs tailored to a specific domain. By focusing on a specific domain, we develop models that are not only more compact in size but also deliver significantly more accurate and contextually relevant responses compared to general-purpose LLMs. Their smaller size enhances efficiency, optimizing both computational resource usage and training time.
@@
Despite the potential of domain-specific LLMs, there is still no systematic study on what makes a good domain-specific LLM. In this work, we consider finance as the domain of interest and aim to address the following research questions:
@@
Given a strong general-purpose LLM (e.g., Llama3-8b-inst), how to effectively adapt it to a target domain (e.g., finance) by post-training? What criteria are desirable for successful adaptation? What are effective training recipes with respect to data and model?
Figure 1: An overview of our finance-specific post-training framework, FinDaP. It comprises four key components: (1) FinCap, the core expected capabilities, including concepts, reasoning, instruction-following and tasks; (2) FinRec, encompassing both data and model strategies to guide domain-adaptive post-training; (3) FinTrain, which curates training texts and prompts based on the data recipe; and (4) FinEval, a comprehensive evaluation framework designed to assess performance on unseen tasks, categorized into similar and novel, general and domain-specific, and reasoning tasks, using both direct-answer and chain-of-thought (CoT) evaluation methods.
Prior studies Bhatia et al. (2024); Xie et al. (2024a) typically adopt a simplified and informal framework (see §2) in that they evaluate only on a set of domain-specific end tasks such as sentiment analysis and NER, and they simply follow standard post-training stages (CPT, IT and/or PA) without considering their impact or optimizing their recipe for domain-adaptive post-training. This simplified approach can misalign with our broader expectations for a domain-expert LLM. A domain-expert LLM should not only excel at such end tasks but also achieve broader capabilities, such as follow task instructions effectively and reason in a way that aligns with domain-specific knowledge, while retaining general capabilities.
We argue that domain-adaptive post-training poses unique challenges compared to pre-training or general post-training. There are multiple factors to be considered: (1) For a particular target domain, it is essential to establish the desirable capabilities that a domain-expert LLM should possess, as these capabilities serve as a guiding framework for the entire adaptation process; (2) The training recipe should be tailored specifically to adapt an already trained LLM (e.g., Llama3-inst) through post-training. This differs from training a model from scratch or from a base pre-trained checkpoint, as it requires careful consideration of catastrophic forgetting and knowledge transfer from the original LLM, which already possesses strong general knowledge and instruction following capabilities. Each of the standard CPT, IT, and PA stages have different impacts and trade-offs with respect to knowledge forgetting and transfer, as do the in-domain, general-domain datasets, and the mixture of them. Moreover, it should also be designed to support the desired capabilities. For example, improving reasoning capability might require more dense supervision than the final answer level correctness score. (3) The desired quantity and quality of training datasets should be carefully balanced: high-quality general-domain data is required to mitigate forgetting, while diverse data and supervision signals are necessary to learn domain knowledge. (4) Finally, the evaluation methods should align with the desired capabilities. Different evaluation techniques may be required for certain capabilities; for example, chain-of-thought (CoT) Wei et al. (2023) reasoning is often necessary to effectively evaluate reasoning tasks.
In this work, we introduce FinDaP (Figure 1), a novel finance-specific framework designed to incorporated all these factors in domain-specific post-training. To our knowledge, none of the prior studies consider all of them to provide a principled guidance on domain-adaptive post-training. FinDaP integrates four key components: (1) FinCap, a set of core capabilities required for the domain expert LLM, including domain concepts, tasks, instruction following and reasoning; (2) FinRec, a training recipe that jointly performs CPT and IT, and subsequently conducts PA, balancing trade-offs across these stages to mitigate forgetting and improve task generalization. It also proposes to use mixture of in-domain and general domain data in the data recipe, alongside a novel preference alignment method for improving reasoning capability that constructs data using the preference signal in reasoning steps, Stepwise Corrective Preference (SCP) and final answer, Final Answer Preference (FAP); (3) FinTrain, a curated set of training datasets implementing FinRec, which carefully balances quality and diversity; and (4) FinEval, a comprehensive evaluation framework covering a wide range of tasks, including reasoning tasks assessed through CoT.
@@
We apply FinDaP on the instruction-tuned Llama3-8b-instruct (LLaMA, 2024). Our best performing recipe yields Llama-Fin that outperforms all considered baselines, including large open models at the 70B scale and proprietary models like GPT-4o, on tasks that are similar (yet unseen) to the training data. Even on novel tasks that were never encountered in training, Llama-Fin remains competitive and consistently outperforms its base model across all identified capabilities.
2 Related Work
Finance LLM Capabilities Recipe Evaluation
Model Recipe Data Recipe
AdaptLLM Concept CPT CPT: Financial text + heuristic QAs constructed from the text Financial + Classification tasks + Direct answer
PIXIU Task IT IT: Financial tasks Financial + Classification tasks + Direct answer
FinLLM Concept, Task CPT → IT CPT: Financial text + Fineweb; IT: Filtered Financial tasks Financial + Classification tasks + Direct answer
FinTral Concept, Task CPT → IT → PA CPT: Financial text; IT: Financial tasks; PA: Outcome signal only Financial + Classification tasks + Direct answer
Palmyra-Fin SoTA public checkpoint, but recipe is not disclosed
Llama-Fin
Concept, IF/Chat,
Task, Reasoning
CPT+IT → PA
CPT: Financial + General text.
IT: Financial + General tasks
PA: A novel PA that leverages outcome and process signals
General + Financial tasks; Similar + Novel tasks
Classification + Open-form QA tasks
Knowledge Recall + Reasoning tasks
Direct answer + CoT
Table 1: Comparison between Llama-Fin with other finance LLMs.
Finance LLMs Table 1 summarizes popular finance-specific LLMs developed through domain-adaptive post-training. AdaptLLM (Cheng et al., 2024) focuses on CPT and constructs heuristic QA tasks from raw text, but it considers only five financial end tasks. PIXIU (Xie et al., 2023) focus on instruction-followingby creating a financial instruction-tuning dataset from diverse open financial tasks and designing a benchmark with nine end tasks for evaluation. FinLLM (Xie et al., 2024a) extends post-training across multiple stages, first performing CPT, then IT, and incorporating multi-modal capabilities via IT. It includes some general-domain data (e.g., fineWeb) but does not explore its impact systematically. Following this line, FinTral (Bhatia et al., 2024) follows this approach but is the only open FinLLM to include PA, where preference labels were given by GPT-4 on the final outcome, considering only coarse-grained signals. It also introduces multi-modality via IT and integrates tool use and retrieval in PA training. Additionally, Palmyra-Fin (Writer, 2024), a recent state-of-the-art FinLLM, reports high performance on finance tasks, particularly CFA exams [^1] https://www.cfainstitute.org/programs/cfa-program/exam, but its training recipe remains undisclosed.
Comparing to FinDaP, none of these models explicitly identify desirable capabilities as we do with FinCap, nor do they systematically explore trade-offs between CPT, IT and PA to develop a more effective training recipe. They also do not incorporate fine-grained process signals in PA to improve reasoning, as we do in FinRec. Additionally, their evaluations lack the broader range of tasks, methods, and similarities, including reasoning tasks and chain-of-thought evaluations, that we adopt in FinEval. Finally, unlike Palmyra-Fin, Llama-Fin is fully open-source, ensuring complete transparency in its training recipe, datasets, and evaluation methods, while achieving SoTA in its size category.
PA for reasoning We explore training-time approaches for improving reasoning (Jiao et al., 2024; DeepSeek-AI et al., 2025). These methods first collect trajectories and then train the LLM with the collected trajectories. This helps the model reason more accurately and faster during inference. To collect reasoning trajectories, there are two main approaches. The first is search-based (Setlur et al., 2024; Snell et al., 2024), where a trained Reward Model (RM) is used to guide a search method (e.g., Best-of-N, Beam Search) to identify the best reasoning path. The second is revision-based (Bai et al., 2022; Du et al., 2023; Madaan et al., 2023; Saunders et al., 2022), which attempts to improve the generation distribution through multi-round interactions, often by leveraging feedback from itself or another strong LLM to refine the input prompt. In practice, revision-based methods have shown mixed results and have not yet been well established as reliable for achieving improvements (Huang et al., 2024a). In contrast, search-based methods have been shown to be more effective. In FinDaP, we propose a novel training-time method that leverages a search-based trajectory collection approach, incorporating both outcome and process rewards from a Generative RM (GenRM).
3 FinDaP Framework
In FinDaP, we first identify four desired capabilities for a finance-expert LLM (FinCap, §3.1). We then develop the training recipe FinRec, which includes both the model recipe that performs CPT and IT jointly followed by PA, and the data recipe, which examines the impact of in-domain, general-domain, and mixed-domain data while introducing a novel data construction approach for PA (§3.2). We then introduce FinTrain3.3), a set of carefully curated training datasets designed to mitigate forgetting while effectively learning domain-specific knowledge. Finally, we propose an evaluation framework FinEval3.4), which considers a diverse set of tasks, ranging from familiar to novel and from general to domain-specific, while also evaluating both direct-answer and CoT methods.
3.1 Core Capabilities (FinCap)
We began by conducting a comprehensive survey of existing work and consulting two financial domain experts—a banking industry advisor and a financial industry product manager. Through this process, we identified four key fundamental capabilities essential for a finance-focused LLM: understanding domain-specific concepts to process financial language accurately, performing domain-specific tasks to solve real-world problems, reasoning effectively to analyze complex financial data, and following instructions to interact naturally in practical applications. These capabilities are deeply interconnected—reasoning depends on conceptual knowledge, while instruction-following ensures effective communication.